OpenStudy (anonymous):

the unit vector of v+3i-j

OpenStudy (anonymous):

v=3i-j

OpenStudy (phi):

divide v by its "length" or magnitude if v = (3,-1) (and it does) |v|^2 = \( v \cdot v \) = 9+1 = 10 and |v|=\(\sqrt{10}\) the unit length vector in the same direction as \(\mathbf{v}\) will be \[ \left(\frac{3}{\sqrt{10}}, -\frac{1}{\sqrt{10}}\right) \] or using i, j notation \[ \frac{3}{\sqrt{10}}\mathbf{i} -\frac{1}{\sqrt{10}}\mathbf{j} \]